Stephanie wants to buy some purses that are on sale for 30% off the original price of $120. The shipping cost is $15. Write an algebraic expression to represent the total cost of the purses. Let x represent the number of purses she is buying.


Answer by ReadingBoosters(3246)

Convert 30% into a decimal
30/100 = .3
Calculate the discount
.3(120) = 36 off the original 120 price
120 - 36 = 84 for each purse
84x+15 = total cost of purses
As one combined expression
[120-(.3)(120)]x + 15 = total cost